>>please write in steps and please explain each.
thank you for your help
Solve for x:
2x3+9x2+4x = 0 <<

Answer by AnlytcPhil(1806)

>>please write in steps and please explain each.
thank you for your help
Solve for x:
2x3+9x2+4x = 0 <<
Factor out x on the left:
x(2x2+9x+4) = 0
Factor the trinomial in parentheses:
x(x+4)(2x+1) = 0
Set each factor = 0
x = 0 (that's one solution)
x+4 = 0
x = -4 (that's a second solution)
2x+1 = 0
2x = -1
x = -1/2 (that's the third solution)
Edwin